OpenStudy (anonymous):

need to simplify. Exponetial notation with positive exponents (9^9)^7

OpenStudy (anonymous):

\[(a^x)^y = a^{xy}\]

OpenStudy (anonymous):

That help?

OpenStudy (anonymous):

(9^9)^7=9^97??

OpenStudy (anonymous):

No, no. when you see "xy" that means "x times y"

OpenStudy (anonymous):

9^56?

OpenStudy (anonymous):

How did you get 56?

OpenStudy (anonymous):

i meant 9^63

OpenStudy (anonymous):

wait, im messing up

OpenStudy (anonymous):

That's correct: \[(9^9)^7 = 9^{9*7} = 9^{63}\]

OpenStudy (anonymous):

i second guessed myself

OpenStudy (anonymous):

so it is 9^63

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

Np